32 oz=_lb, 16lb_oz, 18,000lb=_lb, 21pt=_fl oz
Answer by rfer(16322)

32 oz=2 lbs
16 lbs=16*16=256 oz
18000 lbs=9 tons
21 pt=21*16=336 fl oz